OpenStudy (anonymous):

A six sided die is weighted so that the probabilities of rolling a 1, 2, or 3 are equally likely; the probabilities of rolling a 4, 5, or 6 are equally likely, and the probability of rolling a 5 is four times the probability of rolling a 3. What is the probability that a 2 or a 5 is rolled?

OpenStudy (anonymous):

1=x 2=x 3=x 4=4x 5=4x 6=4x so 15x=1 as sum of probabilities has to equal 1 therefore x=1/15 p(rolling a 2) = x p(rolling a 5)=4*x is that clear?

OpenStudy (anonymous):

Yep, thank you!

OpenStudy (anonymous):

You're welcome

OpenStudy (anonymous):

It also seems intuitive that Prob (1 or 4), Prob(2 or 5) and Prob(3 or 6) would all be equally likely and sum to 1. Thus, each of these has a probability of 1/3